# Instability of Legendrian knottedness, and non-regular Lagrangian concordances of knots

### v1

@@ 1.1. Instability of Legendrian knottedness
The standard Legendrian $\tt{tb}=-1$ unknot $U\subset(\mathbb{R}^{3},dz-y\,dx)$ of the standard contact vector space has zero rotation number and, the standard representative, has a Chekanov–Eliashberg algebra with $\mathbb{Z}[t^{\pm 1}]$-coefficients generated by a single generator $a$ in degree $|a|=1$, and with differential given by $\partial a=1+t$. The first examples of Legendrian knots inside the contact vector-space that are not Legendrian isotopic to $U$, but which have stable-tame isomorphic Chekanov–Eliashberg algebras, were first constructed in [7] by Cornwell–Ng–Sivek. These examples where further explored by Etnyre–Ng in [22, Section 3.4], where it was shown that the Legendrian representatives $\Lambda_{m}$ of the pretzel knots $P(3,-3,-m)$, $m>3$ described in Figure 1 all satisfy this property. Similarly to $U$, the latter Legendrian knots are also Lagrangian slice. However, they are not doubly-slice as shown by Chantraine–Legout [5] and. as a consequence, not Lagrangian cobordant to $U$.
@@
Figure 1. The front projection of the Legendrian representative $\Lambda_{m}$ of the pretzel knot $P(3,-3,m)$ has $m-3$ crossings in the red box.
Recall that the front spinning construction (sometimes called the Legendrian suspension) is a construction that from a given Legendrian submanifold $\Lambda$ in the standard contact vector space $\mathbb{R}^{2n+1}$ produces the Legendrian embedding $\Sigma_{S^{k}}\Lambda$ of $\Lambda\times S^{k}$ in $\mathbb{R}^{2(k+n)+1}$. This construction for $k=1$ was introduced by Ekholm, Etnyre and Sullivan in [15], and then extended to the case of an arbitrary $k$ by the second author in [26]. This construction admits several extra extensions, see [1, 12, 30, 33]. One can also produce a higher-dimensional Legendrian from a lower-dimensional one by taking the Legendrian product of $\Lambda\subset(Y,\alpha)$ with the exact Lagrangian zero-section $0_{M}\subset T^{*}M$. This gives the Legendrian embedding $\Lambda\times 0_{M}\subset(Y\times T^{*}M,\alpha+p\,dq)$. When $M=S^{n}$ and $Y=\mathbb{R}^{3}$ is the contact vector space, the contact manifold $\mathbb{R}^{3}\times T^{*}S^{n}$ admits a canonical embedding into $\mathbb{R}^{3+2n}$. The front-spun $\Sigma_{S^{n}}\Lambda$ can be identified with the image of the Legendrian product under this embedding. Note that both the above Legendrian product, as well as the front-spun, produces a well-defined Legendrian isotopy class, which only depends of the Legendrian isotopy class of the initial Legendrian $\Lambda$.
Here we investigate the Legendrian product operation and front spinning construction when applied to the above family of knots $\Lambda_{m}$. Since these knots only can be distinguished by smooth topological invariants, while they Legendrian invariants agree, it is natural to expect that their products and spuns all become Legendrian isotopic; see the following two remarks.
@@
The Legendrian knots $U$ and $\Lambda_{m}$ all have the same classical invariants except smooth knot class, i.e. the same rotation number and Thurston–Bennequin invariant. Hence, their spuns have the same classical invariants by [15, Lemma 4.16] and [26, Lemma 5.1]. Also their Legendrian contact homology DGAs, $\mathcal{A}(\Lambda_{m})$ and $\mathcal{A}(U)$, are stable tame isomorphic for all $m\geq 3$ with arbitrary coefficients. In particular, using the surgery formula from [16], it follows that the partially wrapped Fukaya categories of the two sectors $\mathbb{C}^{2}_{U}$ and $\mathbb{C}^{2}_{\Lambda_{m}}$ are equivalent, where these sectors are obtained by putting stops at the Legendrians $U,\Lambda_{m}\subset\partial_{\infty}\mathbb{C}^{2}$. Since the partially wrapped Fukaya category of a product of sectors satisfies a Künneth-type formula [23], the partially wrapped Fukaya category of the sector $\mathbb{C}^{2}_{\Lambda_{m}}\times T^{*}S^{n}$ with stop $\Lambda_{m}\times 0_{S^{n}}\subset\partial_{\infty}(\mathbb{C}^{2}\times T^{*}S^{n})$ is thus also equivalent to that of $\mathbb{C}^{2}_{U}\times T^{*}S^{n}$ with stop $U\times 0_{S^{n}}\subset\partial_{\infty}(\mathbb{C}^{2}\times T^{*}S^{n})$ whenever $m\geq 3$. Analogous results for the DGA itself should also follow from a version of Künneths formula for Legendrian contact homology from work in progress by Strakoš [36].
@@
In Section 1.3 we give two related, but different constructions of the Legendrian isotopy needed for proving Theorem 1.3. First, note that $\Lambda_{m}$, $m>3$, can be obtained by an ambient Legendrian surgery $(\Lambda_{0})_{\eta_{m}}$ performed on $\Lambda_{0}$ consisting of two unlinked Legendrian unknots of $\tt{tb}=-1$ along a Legendrian surgery arc $\eta_{m}$ with boundary $\partial\eta_{m}\subset\Lambda_{0}$; see Figures 18 and 17. The corresponding exact Lagrangian cobordism is described in Figure 2.
@@
    The product of the surgery arc $\eta_{m}\times 0_{S^{n}}\subset\mathbb{R}^{3}\times T^{*}S^{n}$ is formally Legendrian isotopic to the spun $\eta_{0}\times 0_{S^{n}}$ of a stabilised Legendrian arc $\eta_{0}$ as shown in Figure 19, for which the Legendrian ambient surgery $(\Lambda_{0})_{\eta_{0}}$ produces the Legendrian standard unknot $U$. Moreover, we need this formal Legendrian isotopy to be supported in the complement of the product $\Lambda_{0}\times 0_{S^{n}}$.
The first strategy uses the fact that the spuns $\eta_{m}\times 0_{S^{n}}$ and $\eta_{0}\times 0_{S^{n}}$ are Legendrian isotopic; this follows from Murphy’s $h$-principle for loose Legendrians [32], which can be applied since spuns of stabilised Legendrians are loose, see [10]. It then follows by construction of Legendrian ambient surgery and the standard Legendrian neighborhood theorem that $\Lambda_{m}\times 0_{S^{n}}$ and $U\times 0_{S^{n}}$ are Legendrian isotopic.
@@ 1.2. Non-regular Lagrangian concordances in the symplectisation of $\mathbb{R}^{3}$
The existence of a flexible regular fillings is a crucial ingredient for constructing the contact isotopy in Theorem 1.3. While there exist plenty of regular fillings whose Weinstein complements are not flexible, we unfortunately do not have a criterion for the regularity of Lagrangians inside Weinstein cobordisms. It is expected that all exact Lagrangian fillings in Weinstein domains are regular, as formulated by Eliashberg–Ganatra–Lazarev in [18, Problem 2.5]. Furthermore, it was proposed there that any Lagrangian cobordism with a negative end which is not loose or stabilised Legendrian is regular. (In [11] the authors showed that this condition must be weakened to, at least, the condition that the negative end is not sub-loose, which means that it is not cobordant to a loose Legendrian.)
Non-regular Lagrangian cobordisms were first constructed by Eliashberg–Murphy in high dimensions [21] and Lin [31] in dimension four. In these cases, the negative ends are stabilised (which means loose in higher dimensions). These Lagrangian cobordisms are non-regular for the basic topological reason that they have empty positive ends; such a regular oriented Lagrangian $L$ inside a Weinstein cobordism $W$ must define a non-zero homology class $[L]\in H_{n}(\overline{W},\partial_{-}\overline{W})$ relative the concave boundary of $\overline{W}$.
@@ 1.3. Lagrangian concordance of Legendrian knots is not a partial order
We provide negative answer to the reasonably old question of whether the Lagrangian concordance relation between Legendrian knots is a partial order [3, 34]. By definition, this relation $\Lambda_{-}\preceq_{conc}\Lambda_{+}$ holds whenever there exists a Lagrangian concordance with negative and positive end $\Lambda_{-}$ and $\Lambda_{+}$, respectively. The relation is clearly reflexive and transitive, and it has been asked whether the relation defines a partial order. The concordances produced in Theorem 1.4 show that the Lagrangian concordance relation is not anti-symmetric, and hence is not a partial order. More precisely, since any Lagrangian slice disc of a Legendrian $\Lambda_{-}$ induces a Lagrangian concordance from the standard Legendrian unknot $U$ to $\Lambda_{-}$, i.e. $U\preceq_{conc}\Lambda_{-},$ and since the Lagrangian concordance relation can be seen to be preserved under stabilisation of the two knots [3], we get:
@@ 2.1. Regular Lagrangian cobordisms
A compact Liouville cobordism is a compact exact symplectic cobordism $(\overline{W},d\lambda)$, whose Liouville vector field $\zeta$ is defined uniquely by the equation $d\lambda(\zeta,\cdot)=\lambda(\cdot)$, and which we require to be transverse to the boundary $\partial\overline{W}$; i.e. the boundary is of contact-type, with contact form given by the restriction $\eta|_{T\partial{\overline{W}}}$. By $(\partial_{-}\overline{W},\alpha_{-})$ (resp. $(\partial_{+}\overline{W},\alpha_{+})$) we denote the concave (resp. convex) contact boundary, i.e. the component along which the Liouville vector field points inwards (resp. outwards). Recall that $\partial_{+}\overline{W}\neq\emptyset$ must hold by Stokes’ theorem. The Liouville cobordism is said to be Weinstein if the Liouville vector field is gradient-like for a Morse function.
@@
±$\partial_{\pm}\overline{L}=\partial\overline{L}\cap\partial_{\pm}\overline{W}\subset(\partial_{\pm}\overline{W},\alpha_{\pm}).$
@@
Before we give the stronger definition, we recall that the union
@@
which, when chosen appropriately, are smooth Weinstein cobordism with concave end $\partial_{-}\overline{W}_{L}=\partial_{-}\overline{W}$, and in which $\overline{L}\subset\overline{W}_{L}$ is regular. Furthermore, the natural Weinstein structure on $D^{*}L$ endows $\overline{W}_{L}$ with a Weinstein structure for which the skeleton of $\overline{W}_{L}$ is entirely contained inside $L$ and hence, in particular, all critical points of the Liouville vector field are contained in $L$. More precisely, in addition to $\zeta$ being tangent to $L$, we can assume that:
@@
Using the language of Weinstein pairs from [17] and Weinstein sectors from [24], the above can be reformulated in the following way. Consider the sector $W_{\Lambda_{+}}$ corresponding to the Weinstein pair $(\overline{W},\Lambda_{+})$, in which $\tilde{L}\coloneqq L\cap W_{\Lambda_{+}}$ is naturally included as a compact Lagrangian with boundary $\partial\tilde{L}$ contained inside the sectorial boundary $\partial W_{\Lambda_{+}}$ of the complete sector. In this setup, $L$ is a special regular Lagrangian iff $\tilde{L}$ can be realised as a subset of the skeleton of $W_{\Lambda_{+}}$ for some homotopic Weinstein structure. In this setting, there is a natural complementary sectorial Weinstein cobordism, which is given as the complement of a neighbourhood of $\tilde{L}$ in the sector.
@@ 2.2. Standard moves for Kirby diagrams
Here we discuss the tools from Kirby calculus that we need to understand and modify the regular cobordisms that we are interested in. We refer to work by Gompf [27] for a good introduction to Kirby calculus in the setting of contact topology. Our Kirby diagrams will consist of a number of attaching 0-spheres inside $(\mathbb{R}^{3}_{xyz},dz-ydx)$ which have vanishing $y$–coordinates. The Legendrian arcs are allowed to enter the handles both from the left and right in the front projection, where the arcs are required to have constant $z$–coordinate. The attaching 1-spheres will typically be depicted in red colour, while Legendrian knots will be black. (The latter can be thought of as stops of a Weinstein sector that is built by attaching the one and two-handles.)
@@
Move K–0: Legendrian isotopy in the complement of the attaching spheres: Both Legendrian knots in the contact manifold, as well as attaching 1-spheres, can be deformed by a Legendrian isotopy that supported in the complement of the attaching 0-spheres. By our convention the attaching 0-spheres have a vanishing $y$–coordinate, which means that any Legendrian arc whose front projection has non-zero slope can pass over the small balls where the attaching 0-spheres live. See Figure 3.
@@
Move K–I: Introducing/cancelling a pair of handles: A Weinstein one- and two-handle that are in cancelling position can be either introduced or removed. This is a standard Kirby move; see [13, Section 5] in work by Ding and Geiges for the version in contact topology. This move is shown in Figure 4.
Figure 4. The Kirby move K–I: The Kirby move that consists of cancelling a pair consisting of a Weinstein 1-handle and a Weinstein 2-handle in standard position corresponds to the following deformation of the front projection.
Move K–II: Handle-sliding parallel strands: Handling sliding parallel strands of a Legendrian over the two-handle which is in cancelling position gives rise to a move shown in Figure 5. This is a standard Kirby move; the version in contact manifolds was described by Ding and Geiges in [13]; and Casals and Murphy in [2].
Figure 5. The Kirby move K–II: The bottom figure shows the effect in the front projection of the Kirby move that consists of handle-sliding $k=2$ parallel Legendrian strands shown on top over its Legendrian two-handle whose attaching sphere is shown in red.
Move K–III: Introducing a double-stabilisation: Taken any number of parallel strands that pass through a single one-handle, one can move the top one to the bottom, introducing a double stabilisation of the same. See [27, Figure 17] for one version, also its reflection in the z-axis is a possible move. See Figure 6.
@@
Move K–IV: Moving strands around the attaching sphere: The bottom strand to the left (resp. right) of the left (resp. right) attaching sphere can be moved to the top of the right (resp. left) side of the same attaching sphere. The strand is given an additional downward cusp-edge in the direction oriented away from the attaching sphere. See Figure 7.

### v2

@@ 1.1. Instability of Legendrian knottedness
In this article we investigate a set of Legendrian knots in $(\mathbb{R}^{3},dz-y\,dx)$ that are not smoothly isotopic, admit exact Lagrangian fillings, and have stable-tame isomorphic Chekanov–Eliashberg algebras. (The latter DGAs are not acyclic in view of the fillability.) Recall that the standard Legendrian $\tt{tb}=-1$ unknot $U$ in the standard contact $3$-dimensional vector space has zero rotation number and a Chekanov–Eliashberg algebra with $\mathbb{Z}[t^{\pm 1}]$-coefficients generated by a single generator $a$ in degree $|a|=1$, and with differential given by $\partial a=1+t$. The first examples of Legendrian knots inside the contact vector-space that are not Legendrian isotopic to $U$, but which have stable-tame isomorphic Chekanov–Eliashberg algebras, were first constructed in [7] by Cornwell–Ng–Sivek. These examples where further explored by Etnyre–Ng in [23, Section 3.4], where it was shown that the Legendrian representatives $\Lambda_{m}$ of the pretzel knots P(3,3,m)P(3,-3,-m), $m>3$ described in Figure 1 all satisfy this property. Similarly to $U$, the latter Legendrian knots are also Lagrangian slice. However, they are not doubly-slice as shown by Chantraine–Legout [5] and, as a consequence, not Lagrangian cobordant to $U$.
@@
Figure 1. The front projection of the Legendrian representative $\Lambda_{m}$ of the pretzel knot P(3,3,m)P(3,-3,m) has m3m-3 crossings in the red box.
Recall that the front spinning construction (sometimes called the Legendrian suspension) is a construction that from a given Legendrian submanifold $\Lambda$ in the standard contact vector space $\mathbb{R}^{2n+1}$ produces the Legendrian embedding $\Sigma_{S^{k}}\Lambda$ of $\Lambda\times S^{k}$ in $\mathbb{R}^{2(k+n)+1}$. This construction for $k=1$ was introduced by Ekholm, Etnyre and Sullivan in [16], and then extended to the case of an arbitrary $k$ by the second author in [27]. This construction admits several generalisations in different directions, see [1, 13, 32, 36]. One can also produce a higher-dimensional Legendrian from a lower-dimensional one by taking the Legendrian product of $\Lambda\subset(Y,\alpha)$ with the exact Lagrangian zero-section $0_{M}\subset T^{*}M$. This gives the Legendrian embedding $\Lambda\times 0_{M}\subset(Y\times T^{*}M,\alpha+p\,dq)$. When $M=S^{n}$ and $Y=\mathbb{R}^{3}$ is the contact vector space, the contact manifold $\mathbb{R}^{3}\times T^{*}S^{n}$ admits a canonical embedding into $\mathbb{R}^{3+2n}$. The front-spun $\Sigma_{S^{n}}\Lambda$ can be identified with the image of the Legendrian product under this embedding. Note that both the above Legendrian product, as well as the front-spun, produces a well-defined Legendrian isotopy class, which only depends of the Legendrian isotopy class of the initial Legendrian $\Lambda$.
Here we investigate the Legendrian product operation and front spinning construction when applied to the above family of knots $\Lambda_{m}$. Since these knots only can be distinguished by smooth topological invariants, while the Legendrian invariants agree, it is natural to suspect that their products and spuns all become Legendrian isotopic; see the following two remarks.
@@
The Legendrian knots $U$ and $\Lambda_{m}$ all have the same classical invariants except smooth knot class, i.e. the same rotation number and Thurston–Bennequin invariant. Hence, their spuns have the same classical invariants by [16, Lemma 4.16] and [27, Lemma 5.1]. Also their Legendrian contact homology DGAs, $\mathcal{A}(\Lambda_{m})$ and $\mathcal{A}(U)$, are stable tame isomorphic for all $m\geq 3$ with arbitrary coefficients. In particular, using the surgery formula from [17], it follows that the partially wrapped Fukaya categories of the two sectors $\mathbb{C}^{2}_{U}$ and $\mathbb{C}^{2}_{\Lambda_{m}}$ are equivalent, where these sectors are obtained by putting stops at the Legendrians $U,\Lambda_{m}\subset\partial_{\infty}\mathbb{C}^{2}$. Since the partially wrapped Fukaya category of a product of sectors satisfies a Künneth-type formula [24], the partially wrapped Fukaya category of the sector $\mathbb{C}^{2}_{\Lambda_{m}}\times T^{*}S^{n}$ with stop $\Lambda_{m}\times 0_{S^{n}}\subset\partial_{\infty}(\mathbb{C}^{2}\times T^{*}S^{n})$ is thus also equivalent to that of $\mathbb{C}^{2}_{U}\times T^{*}S^{n}$ with stop $U\times 0_{S^{n}}\subset\partial_{\infty}(\mathbb{C}^{2}\times T^{*}S^{n})$ whenever $m\geq 3$. Analogous results for the DGA itself should also follow from a version of Künneths formula for Legendrian contact homology from work in progress by Strakoš [40].
@@
In Section 4 we give two related, but different constructions of the Legendrian isotopy needed for proving Theorem 1.3. First, note that $\Lambda_{m}$, $m>3$, can be obtained by an ambient Legendrian surgery $(\Lambda_{0})_{\eta_{m}}$ performed on $\Lambda_{0}$ consisting of two unlinked Legendrian unknots of $\tt{tb}=-1$ along a Legendrian surgery arc $\eta_{m}$ with boundary $\partial\eta_{m}\subset\Lambda_{0}$; see Figures 18 and 17. The corresponding exact Lagrangian cobordism is described in Figure 2.
@@
    The product of the surgery arc $\eta_{m}\times 0_{S^{n}}\subset\mathbb{R}^{3}\times T^{*}S^{n}$ is formally Legendrian isotopic to the spun $\eta^{0}_{m}\times 0_{S^{n}}$ of a stabilised Legendrian arc $\eta^{0}_{m}$ as shown in Figure 19, for which the Legendrian ambient surgery $(\Lambda_{0})_{\eta^{0}_{m}}$ produces the Legendrian standard unknot $U$. Moreover, we need this formal Legendrian isotopy to be supported in the complement of the product $\Lambda_{0}\times 0_{S^{n}}$.
The first strategy uses the fact that the spuns $\eta_{m}\times 0_{S^{n}}$ and $\eta^{0}_{m}\times 0_{S^{n}}$ are Legendrian isotopic by an isotopy supported in the complement of $\Lambda_{0}\times 0_{S^{n}}$; this follows from Murphy’s $h$-principle for loose Legendrians [34], which can be applied since spuns of stabilised Legendrians are loose, see [11]. It then follows by construction of Legendrian ambient surgery and the standard Legendrian neighborhood theorem that $\Lambda_{m}\times 0_{S^{n}}$ and $U\times 0_{S^{n}}$ are Legendrian isotopic.
@@ 1.2. Non-regular Lagrangian concordances in the symplectisation of $\mathbb{R}^{3}$
The existence of a flexible regular filling is a crucial ingredient for constructing the contact isotopy in Theorem 1.3. While there exist plenty of regular fillings whose Weinstein complements are not flexible, we unfortunately do not have a good understanding of when regularity holds for Lagrangians inside Weinstein cobordisms. It is expected that all exact Lagrangian fillings in Weinstein domains are regular, as formulated by Eliashberg–Ganatra–Lazarev in [19, Problem 2.5]. Furthermore, it was proposed there that any Lagrangian cobordism with a negative end which is not loose or stabilised Legendrian is regular. (In [12] the authors showed that this condition must be weakened to, at least, the condition that the negative end is not sub-loose, which means that it is not cobordant to a loose Legendrian.)
Non-regular Lagrangian cobordisms were first constructed by Eliashberg–Murphy in high dimensions [22] and Lin [33] in dimension four. In these cases, the negative ends are stabilised (which means loose in higher dimensions). These Lagrangian cobordisms are non-regular for the basic topological reason that they have empty positive ends; a regular oriented Lagrangian $L$ inside a Weinstein cobordism $W$ must define a non-zero homology class $L]\in H_{n}(\overline{W},\partial_{-}\overline{W})$ relative the concave boundary of $\overline{W}$.
@@ 1.3. Lagrangian concordance of Legendrian knots is not a partial order
We provide negative answer to the reasonably old question of whether the Lagrangian concordance relation between Legendrian isotopy classes of Legendrian knots is a partial order [3, 37]. By definition, this relation $\Lambda_{-}\preceq_{conc}\Lambda_{+}$ holds whenever there exists a Lagrangian concordance with negative and positive end $\Lambda_{-}$ and $\Lambda_{+}$, respectively. The relation is clearly reflexive and transitive, and it has been asked whether the relation defines a partial order. The concordances produced in Theorem 1.4 show that the Lagrangian concordance relation is not anti-symmetric, and hence is not a partial order. More precisely, since any Lagrangian slice disc of a Legendrian $\Lambda_{-}$ induces a Lagrangian concordance from the standard Legendrian unknot $U$ to $\Lambda_{-}$, i.e. $U\preceq_{conc}\Lambda_{-},$ and since the Lagrangian concordance relation can be seen to be preserved under stabilisation of the two knots [3], we get:
@@ 2.1. Regular Lagrangian cobordisms
A compact Liouville cobordism is a compact exact symplectic cobordism $(\overline{W},d\lambda)$, whose Liouville vector field $\zeta$ is defined uniquely by the equation $d\lambda(\zeta,\cdot)=\lambda(\cdot)$, and which we require to be transverse to the boundary $\partial\overline{W}$; i.e. the boundary is of contact-type, with contact form given by the restriction $\lambda|_{T\partial{\overline{W}}}$. By $(\partial_{-}\overline{W},\alpha_{-})$ (resp. $(\partial_{+}\overline{W},\alpha_{+})$) we denote the concave (resp. convex) contact boundary, i.e. the component along which the Liouville vector field points inwards (resp. outwards). Recall that $\partial_{+}\overline{W}\neq\emptyset$ must hold by Stokes’ theorem. The Liouville cobordism is said to be Weinstein if the Liouville vector field is gradient-like for a Morse function.
@@
$\partial_{\pm}\overline{L}=\partial\overline{L}\cap\partial_{\pm}\overline{W}\subset(\partial_{\pm}\overline{W},\alpha_{\pm}).$
@@
Before we give the stronger definition, we recall, following the discussion in [19, Section 2], that the union
@@
which, when chosen appropriately, is a smooth compact Weinstein cobordism with concave end $\partial_{-}\overline{W}_{L}=\partial_{-}\overline{W}$, and in which $\overline{L}\subset\overline{W}_{L}$ is regular. Furthermore, the natural Weinstein structure on $D^{*}L$ endows $\overline{W}_{L}$ with a Weinstein structure for which the skeleton of $\overline{W}_{L}$ is entirely contained inside $L$ and hence, in particular, all critical points of the Liouville vector field are contained in $L$. More precisely, in addition to $\zeta$ being tangent to $L$, we can assume that:
@@
Using the language of Weinstein pairs from [18] and Weinstein sectors from [25], the above can be reformulated in the following way. Consider the sector $W_{\Lambda_{+}}$ corresponding to the Weinstein pair $(\overline{W},\Lambda_{+})$, in which $\tilde{L}\coloneqq L\cap W_{\Lambda_{+}}$ is naturally included as a compact Lagrangian with boundary $\partial\tilde{L}$ contained inside the sectorial boundary $\partial W_{\Lambda_{+}}$ of the complete sector. In this setup, $L$ is a special regular Lagrangian if and only if $\tilde{L}$ can be realised as a subset of the skeleton of $W_{\Lambda_{+}}$ for some homotopic Weinstein structure. In this setting, there is a natural complementary sectorial Weinstein cobordism, which is given as the complement of a neighbourhood of $\tilde{L}$ in the sector.
@@ 2.2. Standard moves for Kirby diagrams
Here we discuss the tools from Kirby calculus that we need to understand and modify the regular cobordisms that we are interested in. We refer to work by Gompf [28] for a good introduction to Kirby calculus in the setting of contact topology. Our Kirby diagrams will consist of a number of attaching 0-spheres inside $(\mathbb{R}^{3}_{xyz},dz-ydx)$ which have vanishing yy–coordinates. The Legendrian arcs are allowed to enter the handles both from the left and right in the front projection, where the arcs are required to have constant zz–coordinate. The attaching 1-spheres will typically be depicted in red colour, while Legendrian knots will be black. (The latter can be thought of as stops of a Weinstein sector that is built by attaching the one and two-handles.)
@@
Move K–0: Legendrian isotopy in the complement of the attaching spheres: Both Legendrian knots in the contact manifold, as well as attaching 1-spheres, can be deformed by a Legendrian isotopy that supported in the complement of the attaching 0-spheres. By our convention the attaching 0-spheres have a vanishing yy–coordinate, which means that any Legendrian arc whose front projection has non-zero slope can pass over the small balls where the attaching 0-spheres live. See Figure 3.
@@
Move K–I: Introducing/cancelling a pair of handles: A Weinstein one- and two-handle that are in cancelling position can be either introduced or removed. This is a standard Kirby move; see [14, Section 5] in work by Ding and Geiges for the version in contact topology. This move is shown in Figure 4.
Figure 4. The Kirby move K–I: The Kirby move that consists of cancelling a pair consisting of a Weinstein 1-handle and a Weinstein 2-handle in standard position corresponds to the above deformation of the front projection.
Move K–II: Handle-sliding parallel strands: Handling sliding parallel strands of a Legendrian over a Weinstein two-handle which is in cancelling position gives rise to a move shown in Figure 5. This is a standard Kirby move; the version in contact manifolds was described by Ding and Geiges in [14]; and Casals and Murphy in [2].
Figure 5. The Kirby move K–II: The bottom figure shows the effect in the front projection of the Kirby move that consists of handle-sliding $k=2$ parallel Legendrian strands shown on top over the Weinstein two-handle whose attaching sphere is the knot shown in red.
Move K–III: Introducing a double-stabilisation: Given any number of parallel strands that pass through a single one-handle, one can move the bottom strand to the top, introducing a double stabilisation of the same. See [28, Figure 17] for one version, also its reflection in the z-axis is a possible move. See Figure 6.
@@
Move K–IV: Moving strands around the attaching sphere: The bottom strand to the left (resp. right) of the left (resp. right) attaching sphere can be moved to the top of the right (resp. left) side of the same attaching sphere. The strand is given an additional downward cusp-edge in the direction oriented away from the attaching sphere. See Figure 7 for the move in the front projection, and Figure 8 for how the move can be described using the Lagrangian projection.